OpenStudy (reedplymel):

Find the difference: (6y3 + 17y − 3) − (4y3 − 11y + 9) 2y3 + 28y + 12 2y3 + 6y − 12 2y3 + 28y − 6 2y3 + 28y − 12

OpenStudy (welshfella):

first distribute the negative over the second parentheses

OpenStudy (anonymous):

D

OpenStudy (reedplymel):

the last one??

OpenStudy (anonymous):

Yes its the last one

OpenStudy (reedplymel):

ok medal goes to u thx

OpenStudy (anonymous):

yay thank you

OpenStudy (reedplymel):

Find the sum: (3x2 + 5x − 8) + (5x2 − 13x − 5) 8x2 + 8x − 13 8x2 − 8x + 13 8x2 − 8x − 13 8x2 − x − 13

OpenStudy (reedplymel):

this one

OpenStudy (anonymous):

c

OpenStudy (reedplymel):

ok can u work with me on the rest so i can know tjis stuff instead of u giving it to mee Simplify the expression (2x − 9)(x + 6). 2x2 − 54 2x2 − 3x − 54 2x2 + 3x − 54 2x2 + 21x − 54

OpenStudy (welshfella):

@beckyg111 Please dont just give the answer - guide the questioner through the problem

OpenStudy (reedplymel):

ya im tryin to learn can u help me

OpenStudy (anonymous):

Ok sorry @welshfella can u help him i need to take a test

OpenStudy (welshfella):

(2x − 9)(x + 6) first you multiply the second part by 2x then multiply it by -9 = 2x(x + 6) - 9(x + 6) ok so far?

OpenStudy (reedplymel):

yeah

OpenStudy (welshfella):

right- now you distribute the 2x and the -9 over the 2 parentheses = 2x*x + 2x*6 - 9*x -9*6 = 2x^2 + 12x - 9x - 54 notice the last term is negative 54 because you multiply negative times positive

OpenStudy (reedplymel):

so A is out?

OpenStudy (welshfella):

now the 2 middle terms are alike so it simplifies to 2x^2 + 3x - 54

OpenStudy (welshfella):

yes its not A

OpenStudy (reedplymel):

ok were did u get the 3x

OpenStudy (welshfella):

12x - 9x = 3x

OpenStudy (welshfella):

- combining like terms

OpenStudy (reedplymel):

theres only one x plus the 2x aint no 12

OpenStudy (welshfella):

2x^2 + 12x - 9x - 54 = 2x^2 + 3x - 54 the 2 middle terms are 12x and - 9x - they are like termsa so you can simplify them

OpenStudy (reedplymel):

ohh so its C?

OpenStudy (welshfella):

yes

OpenStudy (reedplymel):

mb computer messin up

OpenStudy (reedplymel):

thx

OpenStudy (welshfella):

yw

OpenStudy (reedplymel):

do u think u can do some more wit me

OpenStudy (welshfella):

i can probably do one more I have to go soon.

OpenStudy (reedplymel):

ya me too Simplify the expression -2x2(4x − 3).

OpenStudy (reedplymel):

8x3 − 6x2 -8x3 + 6x2 8x3 − 6x2 8x3 + 6x2

OpenStudy (welshfella):

= -2x^2 * 4x - 2x^2 * -3

OpenStudy (welshfella):

-2 * 4 = ?

OpenStudy (welshfella):

remember negative * positive = negative